# UniEmoX: Cross-modal Semantic-Guided Large-Scale Pretraining for Universal Scene Emotion Perception

### v2

@@ Abstract
Visual emotion analysis holds significant research value in both computer vision and psychology. However, existing methods for visual emotion analysis suffer from limited generalizability due to the ambiguity of emotion perception and the diversity of data scenarios. To tackle this issue, we introduce UniEmoX, a cross-modal semantic-guided large-scale pretraining framework. Inspired by psychological research emphasizing the inseparability of the emotional exploration process from the interaction between individuals and their environment, UniEmoX integrates scene-centric and person-centric low-level image spatial structural information, aiming to derive more nuanced and discriminative emotional representations. By exploiting the similarity between paired and unpaired image-text samples, UniEmoX distills rich semantic knowledge from the CLIP model to enhance emotional embedding representations more effectively. To the best of our knowledge, this is the first large-scale pretraining framework that integrates psychological theories with contemporary contrastive learning and masked image modeling techniques for emotion analysis across diverse scenarios. Additionally, we develop a visual emotional dataset titled Emo8. Emo8 samples cover a range of domains, including cartoon, natural, realistic, science fiction and advertising cover styles, covering nearly all common emotional scenes. Comprehensive experiments conducted on six benchmark datasets across two downstream tasks validate the effectiveness of UniEmoX. The source code is available at https://github.com/chincharles/u-emo.
@@
publicationid: pubid: $0$–$00$ 2021 IEEE
Figure 1: Overview of UniEmoX: A cross-modal semantic-guided large-scale emotion pretraining framework for downstream emotion analysis tasks.
@@ I Introduction
In recent years, large-scale vision models [11, 12] have achieved tremendous success, with their performance not yet significantly limited by the capacity of the model. Inspired by the paradigm of pre-training large-scale vision models, we can learn rich semantic and visual representations from large-scale datasets. Subsequently, fine-tuning on corresponding downstream datasets helps address the issue of lack of generalizability in visual emotion analysis methods. We ask: How to learn general emotional representations from large-scale complex scenes for effective application in various downstream emotion analysis tasks?
@@
In summary, we propose UniEmoX, a cross-semantic guided multimodal pretraining framework. Our aim is to integrate popular pretraining techniques, such as masked image modeling and contrastive learning methods, with psychologically validated emotion-related theory prior knowledge to construct a universal emotion representation framework. Although our approach may appear as a straightforward integration of existing techniques, the core challenge is ensuring that this integration outperforms the use of individual methods in isolation. These difficulties can be summarized in three key aspects. First, most existing visual emotion datasets lack a textual modality, making it challenging to directly apply CLIP, a multimodal model, in this context. Building a new visual emotion dataset with textual descriptions from scratch would require substantial time and effort. Second, determining how to appropriately incorporate psychological prior knowledge—particularly the theory that emotions are inseparable from interactions between humans and their environment—remains a pressing issue. Finally, while masked image modeling and contrastive learning techniques are relatively mature individually, ensuring stability during training and optimizing performance when combining these two methods presents another critical challenge.
To address the first difficulty, We construct 118,000 pairs of emotional images and their corresponding textual descriptions from the EmoSet dataset, in which the descriptions are organized using emotion attribute-based natural language logic. Second, we develop a similarity-based contrastive loss method to effectively distill visual semantic knowledge from CLIP. To tackle the second difficulty, we propose four fusion strategies designed to bridge the various semantic gaps between humans and their environment. Finally, in response to the third difficulty, we introduce a dual-stream multi-loss joint optimization framework grounded in a masked autoencoder, incorporating various techniques during the training process, including data augmentation and continuous pre-training strategies.
@@
  • We develop UniEmoX, a pretraining framework for emotion analysis in universal scenes. UniEmoX effectively addresses the limited generalizability that current emotion analysis methods encounter due to the ambiguity of emotion perception and the diversity of data scenarios.
  • In the development of UniEmoX, we address three primary difficulties. First, to overcome the limitations of textual modalities and the design of contrastive framework, we integrate emotional attributes into texts through natural language logic, constructing 118,000 emotion image-text pairs from the EmoSet dataset. We devise a similarity-based contrastive loss method to effectively extract visual semantic knowledge from CLIP. Second, To apply psychological theories practically and bridge the semantic gap between humans and their environment, we propose four fusion strategies. Lastly, to improve performance and training stability, we establish a dual-stream, multi-loss joint optimization framework based on masked autoencoding, incorporating various techniques such as data augmentation and continuous pretraining throughout the training process.
@@
  • We conduct extensive experiments on two downstream tasks: scene-centric visual emotion analysis and person-centric visual emotion analysis, as well as on six benchmark datasets to validate the effectiveness of our proposed UniEmoX.
@@ II Related Work
In this section, we review existing works related to our UniEmoX from scene-centric visual emotion analysis, person-centric visual emotion analysis, and large-scale visual model pre-training.
@@ II-B Person-centric Visual Emotion Analysis
Person-centered visual emotion analysis (PVEA) aims to determine the emotional state of specific targets by considering both background and target information. Current methods can be broadly classified into two categories. The first category primarily focuses on exploring various emotion-related features to design sophisticated network structures capable of effectively extracting and integrating these features. Early study [7] highlights the significance of contextual information, which leads to the development of the EmotiCon [8]. EmotiCon introduces types of contexts (gait, facial features, background, and social dynamics) from a psychological perspective [23, 24, 25, 7] to enrich emotional features. Subsequent study [26] developes diverse mechanisms for extracting both foreground and background features, integrating multi-dimensional emotional information to enhance the model’s expressive capabilities. The second category focuses on improving classification performance by addressing data background bias and developing innovative multi-task learning strategies. To mitigate background bias, a post-hoc adjustment method known as the Context Causal Intervention Module (CCIM) [27] proposes to counteract the influence of background information on classification outcomes. Additionally, the multi-adaptive optimization (MAO) strategy [28] introduces an adaptive optimization technique that dynamically adjusts the contribution of each task to the model parameters, thereby enhancing overall classification performance. However, despite the strong performance of these methods on specific datasets, they, similar to SVEA methods, face challenges in achieving widespread deployment in complex scenarios due to their intricate structures and limited generalization capabilities.
@@ II-C Large-Scale Visual Model Pre-training
In computer vision, the pre-training paradigm can be divided into two main types: supervised pre-training [13] and unsupervised pre-training. Supervised pre-training involves training on large-scale labeled datasets, followed by fine-tuning for specific tasks. This study focuses on self-supervised pre-training, a form of unsupervised pre-training that does not rely on labeled data. Instead, it extracts supervisory information from large-scale unlabeled data by designing auxiliary tasks, known as pretext tasks. Self-supervised pre-training methods can be categorized into several types based on different design philosophies: data augmentation-based methods, contrastive learning-based methods, and masked image modeling-based methods. Data augmentation-based methods utilize inherent contextual relationships between samples, such as spatial structure and the maintenance of local and global consistency, to construct supervisory signals. For example, Researchers [16] train DNNs to learn image representations by recognizing random geometric transformations. Additionally, tasks like color prediction [17, 29] and jigsaw puzzles [18] have been employed as pre-training tasks for self-supervised learning. Contrastive learning-based methods aim to maximize the similarity between positive samples while minimizing the similarity between negative samples. Starting from the simple instance discrimination task, research has evolved to include MoCoV3 [30], DINO [31] and dBOT [32], up to the more popular CLIP [12], showcasing the superiority of contrastive learning. Masked image modeling-based methods leverage the co-occurrence relationships between image patches as supervisory signals. For example, BEiT [33] uses the token output of a pre-trained tokenizer as its target, while MAE [11] directly uses the raw pixels as targets. However, the primary aim of these works is to learn low-level visual patterns. Our goal is to learn abstract visual emotion representations by integrating popular pretraining techniques, such as masked image modeling and contrastive learning, with psychology-validated theories of emotion. We address the challenges of combining theoretical methods and establish a versatile and effective emotion representation framework, UniEmoX. UniEmoX derives rich scene emotion semantics from the large-scale EmoSet dataset and improves generalization performance and deployment by substituting complex network designs with parameter-detachable ViT structures.
@@
Figure 5: Network architecture of the cross-semantic guided multimodal pretraining framework (UniEmoX).
@@ III Methodology
In this section, we provide an overview of the proposed UniEmoX, followed by a detailed discussion of its network architecture. The development of UniEmoX is influenced by popular pretraining techniques, such as masked image modeling and contrastive learning, as well as psychological theories of emotion. These influences contribute to methods for human-environment (scene) image reconstruction, the distillation of rich semantic knowledge from CLIP, and discriminative emotional representation learning. Finally, we discuss the optimization objectives during the fine-tuning stage of downstream tasks.
@@ III-A Overview
III-B Network Architecture
Human-Environment (Scene) Image Reconstruction. Inspired by Masked Autoencoders (MAE), we enhance the model’s capability to automatically capture and learn implicit structures and contextual information within images by reconstructing masked regions during pre-training phase. This approach improves the model’s perception of both image details and global semantics.
@@ III-B Network Architecture
We utilize the design of the loss function from MAE to encourage the model to learn meaningful data representation by minimizing the difference between the reconstructed image sequence $S^{\prime}$ of the masked parts and the original image sequence $S$. The loss function $\mathcal{L}_{1}$ is expressed as follows
@@
Discriminative Emotional Representation Learning. According to psychological prior theories, the cognitive process of emotion is closely tied to the interaction between an individual and their environment. However, a natural semantic gap exists between these theories and the design of deep learning models. A significant challenge lies in designing effective fusion modules that integrate the feature representations of the full-scene image $I$ and the human position map $B$ to derive more nuanced and discriminative emotional representations. To address this challenge, we develop four methods, each targeting different degrees of this semantic gap.
@@
Before introducing these four fusion methods, we first explain the process of feature representation for human position map. The full-scene image $I$ is processed by a YOLOv3 detector [35] and undergoes an image scaling transformation to obtain the human position map $B\in\mathbb{R}^{N\times H\times W\times C_{1}}$. The human position map $B$ is subjected to similar operations as $I$, passing through the ViT encoder and a linear layer to extract the CLS-token, denoted as $\beta\in\mathbb{R}^{N\times 1\times C_{3}}$. The feature representations of the body position map $\beta$ and the full-scene image $\alpha$ are reduced in dimensionality to $\mathbb{R}^{N\times C_{3}}$ and $\mathbb{R}^{N\times C_{3}}$, respectively.
To fuse features $\alpha$ and $\beta$ from human position map $B$ and full-scene image $I$, we propose four methods to address varying levels of semantic gaps. To effectively tackle strong semantic gaps, we design fusion strategies $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$, which utilize weighted sum fusion, multi-head attention mechanisms with advanced activation functions, and complex feature mapping fusion, respectively. Additionally, as shown in the Psychological Prior section in the upper right corner of Fig. 5, we develop a simple feature addition strategy as a baseline for comparison to evaluate the effectiveness of the other fusion strategies in bridging semantic gaps.
For the fusion strategy $\Gamma_{1}$, multiple sets of parameters $\lambda_{i}\in\mathbb{R}^{1\times C_{3}}$ and $\mu_{i}\in\mathbb{R}^{1\times C_{3}}$ are defined for $\forall i\in\left\{1,2,\dots,\kappa\right\}$. In $\Gamma_{1}$, the weighted sum of features $\alpha$ and $\beta$ from each set $\lambda_{i}$ and $\mu_{i}$ is combined and then subjected to layer normalization. To prevent issues with zero values during computation, a small bias term $\epsilon\in\mathbb{R}^{C_{3}}$ is added. The fusion strategy $\Gamma_{1}$ is expressed as follows
@@
$\Upsilon=LayerNorm\left(\sum_{i=1}^{\kappa}\left(\lambda_{i}\odot\alpha+\mu_{i}\odot\beta\right)\right)+\epsilon.$ (2)

### v3

@@ Abstract
Visual emotion analysis holds significant research value in both computer vision and psychology. However, existing methods for visual emotion analysis suffer from limited generalizability due to the ambiguity of emotion perception and the diversity of data scenarios. To tackle this issue, we introduce UniEmoX, a cross-modal semantic-guided large-scale pretraining framework. Inspired by psychological research emphasizing the inseparability of the emotional exploration process from the interaction between individuals and their environment, UniEmoX integrates scene-centric and person-centric low-level image spatial structural information, aiming to derive more nuanced and discriminative emotional representations. By exploiting the similarity between paired and unpaired image-text samples, UniEmoX distills rich semantic knowledge from the CLIP model to enhance emotional embedding representations more effectively. To the best of our knowledge, this is the first large-scale pretraining framework that integrates psychological theories with contemporary contrastive learning and masked image modeling techniques for emotion analysis across diverse scenarios. Additionally, we develop a visual emotional dataset titled Emo8. Emo8 samples cover a range of domains, including cartoon, natural, realistic, science fiction and advertising cover styles, covering nearly all common emotional scenes. Comprehensive experiments conducted on seven benchmark datasets across two downstream tasks validate the effectiveness of UniEmoX. The source code is available at https://github.com/chincharles/u-emo.
@@
Figure 1: Overview of UniEmoX: A cross-modal semantic-guided large-scale emotion pretraining framework for downstream visual emotion analysis tasks.
@@ I Introduction
In recent years, large-scale vision models [11, 12] have achieved tremendous success, with their performance not yet significantly limited by the capacity of the model. Inspired by the paradigm of pre-training large-scale vision models, we can learn rich semantic and visual representations from large-scale datasets. Subsequently, fine-tuning on corresponding downstream datasets helps address the issue of lack of generalizability in visual emotion analysis methods. We ask: How to learn general emotional representations from large-scale complex scenes for effective application in various downstream visual emotion analysis tasks?
@@
In summary, we propose UniEmoX, a cross-semantic guided multimodal pretraining framework. Our aim is to integrate popular pretraining techniques, such as masked image modeling and contrastive learning methods, with psychologically validated emotion-related theory prior knowledge to construct a universal emotion representation framework. Although our approach may appear as a straightforward integration of existing techniques, the core challenge is ensuring that this integration outperforms the use of individual methods in isolation. These difficulties can be summarized in three key aspects. First, most existing visual emotion datasets lack a textual modality, making it challenging to directly apply CLIP, a multimodal model, in this context. Building a new visual emotion dataset with textual descriptions from scratch would require substantial time and effort. Efficiently transferring visual-semantic knowledge from CLIP to an emotion pretraining framework remains a challenging problem. Second, determining how to appropriately incorporate psychological prior knowledge—particularly the theory that emotions are inseparable from interactions between humans and their environment—remains a pressing issue. Finally, while masked image modeling and contrastive learning techniques are relatively mature individually, ensuring stability during training and optimizing performance when combining these two methods presents another critical challenge.
To address the first difficulty, we construct 118,000 pairs of emotional images and their corresponding textual descriptions from the EmoSet dataset, in which the descriptions are organized using emotion attribute-based natural language logic. Second, we propose a Contrastive Feature Alignment (CFA) loss and a Semantic-aware Visual Alignment (SVA) loss to effectively distill general visual-semantic knowledge from CLIP. To tackle the second difficulty, we propose four fusion strategies designed to bridge the various semantic gaps between humans and their environment. Finally, in response to the third difficulty, we introduce a dual-stream multi-loss joint optimization framework grounded in a masked autoencoder, incorporating various techniques during the training process, including data augmentation and continuous pre-training strategies.
@@
  • We develop UniEmoX, a pretraining framework for emotion analysis in universal scenes. UniEmoX effectively addresses the limited generalizability that current visual emotion analysis methods encounter due to the ambiguity of emotion perception and the diversity of data scenarios.
  • In the development of UniEmoX, we address three primary difficulties. First, to overcome the limitations of textual modalities and the design of contrastive framework, we integrate emotional attributes into texts through natural language logic, constructing 118,000 emotion image-text pairs from the EmoSet dataset. We propose a Contrastive Feature Alignment (CFA) loss and a Semantic-aware Visual Alignment (SVA) loss to effectively distill general visual-semantic knowledge from CLIP. Second, To apply psychological theories practically and bridge the semantic gap between humans and their environment, we propose four fusion strategies. Lastly, to improve performance and training stability, we establish a dual-stream, multi-loss joint optimization framework based on masked autoencoding, incorporating various techniques such as data augmentation and continuous pretraining throughout the training process.
@@
  • We conduct extensive experiments on two downstream tasks: scene-centric visual emotion analysis and person-centric visual emotion analysis, as well as on seven benchmark datasets to validate the effectiveness of our proposed UniEmoX.
@@ II Related Work
In this section, we review existing works related to our UniEmoX from scene-centric visual emotion analysis, person-centric visual emotion analysis, and large-scale visual model pre-training. Further details are available in review articles [20, 21] on visual emotion analysis.
@@ II-B Person-centric Visual Emotion Analysis
Person-centered visual emotion analysis (PVEA) aims to determine the emotional state of specific targets by considering both background and target information. Current methods can be broadly classified into two categories. The first category primarily focuses on exploring various emotion-related features to design sophisticated network structures capable of effectively extracting and integrating these features. Early study [7] highlights the significance of contextual information, which leads to the development of the EmotiCon [8]. EmotiCon introduces types of contexts (gait, facial features, background, and social dynamics) from a psychological perspective [25, 26, 27, 7, 28, 29] to enrich emotional features. Subsequent study [30] developes diverse mechanisms for extracting both foreground and background features, integrating multi-dimensional emotional information to enhance the model’s expressive capabilities. The second category focuses on improving classification performance by addressing data background bias and developing innovative multi-task learning strategies. To mitigate background bias, a post-hoc adjustment method known as the Context Causal Intervention Module (CCIM) [31] proposes to counteract the influence of background information on classification outcomes. Additionally, the multi-adaptive optimization (MAO) strategy [32] introduces an adaptive optimization technique that dynamically adjusts the contribution of each task to the model parameters, thereby enhancing overall classification performance. However, despite the strong performance of these methods on specific datasets, they, similar to SVEA methods, face challenges in achieving widespread deployment in complex scenarios due to their intricate structures and limited generalization capabilities.
@@ II-C Large-Scale Visual Model Pre-training
In computer vision, the pre-training paradigm can be divided into two main types: supervised pre-training [13] and unsupervised pre-training. Supervised pre-training involves training on large-scale labeled datasets, followed by fine-tuning for specific tasks. This study focuses on self-supervised pre-training, a form of unsupervised pre-training that does not rely on labeled data. Instead, it extracts supervisory information from large-scale unlabeled data by designing auxiliary tasks, known as pretext tasks. Self-supervised pre-training methods can be categorized into several types based on different design philosophies: data augmentation-based methods, contrastive learning-based methods, and masked image modeling-based methods. Data augmentation-based methods utilize inherent contextual relationships between samples, such as spatial structure and the maintenance of local and global consistency, to construct supervisory signals. For example, Researchers [16] train DNNs to learn image representations by recognizing random geometric transformations. Additionally, tasks like color prediction [17, 33] and jigsaw puzzles [18] have been employed as pre-training tasks for self-supervised learning. Contrastive learning-based methods aim to maximize the similarity between positive samples while minimizing the similarity between negative samples. Starting from the simple instance discrimination task, research has evolved to include MoCoV3 [34], DINO [35] and dBOT [36], up to the more popular CLIP [12], showcasing the superiority of contrastive learning. Masked image modeling-based methods leverage the co-occurrence relationships between image patches as supervisory signals. For example, BEiT [37] uses the token output of a pre-trained tokenizer as its target, while MAE [11] directly uses the raw pixels as targets. However, the primary aim of these works is to learn low-level visual patterns. Our goal is to learn general visual emotion representations by integrating popular pretraining techniques, such as masked image modeling and contrastive learning, with psychology-validated theories of emotion. We address the challenges of combining theoretical methods and establish a versatile and effective emotion representation framework, UniEmoX. UniEmoX derives rich scene emotion semantics from the large-scale EmoSet dataset and improves generalization performance and deployment by substituting complex network designs with parameter-detachable ViT structures.
@@
Figure 5: Network architecture of the cross-modal semantic-guided large-scale emotion pretraining framework.
@@ III Methodology
We propose UniEmoX, an emotion pretraining framework that synergistically integrates psychological prior knowledge with self-supervised learning methods to address the limitations in current visual emotion analysis methods. Section III-A first introduces the foundational architecture and core design principles. The architecture comprises three innovative modules: (1) Human-Scene Image Reconstruction (Section III-B) explicitly models contextual interaction features between individuals and scenes; (2) Discriminative Emotional Representation Learning (Section III-C) bridges cross-modal semantic gaps through hierarchical feature fusion; and (3) Distilling Rich Semantic Knowledge from CLIP (Section III-D) transfers emotion-relevant knowledge from vision-language models. Section III-E optimizes task-specific generalization and cross-domain adaptability.
@@ III-A Overview
III-B Human-Scene Image Reconstruction
Inspired by Masked Autoencoders (MAE), we enhance the model’s capability to automatically capture and learn implicit structures and contextual information of both individuals and their environments by reconstructing masked regions during pre-training phase. This approach improves the model’s perception of both the details and global semantics of individuals and their surroundings.
@@ III-B Human-Scene Image Reconstruction
We utilize the design of the loss function from MAE to encourage the model to learn meaningful data representation by minimizing the difference between the reconstructed image sequence $S^{\prime}$ of the masked parts and the original image sequence $S$. The loss function $\mathcal{L}_{1}$ is defined as follows:
@@
As presented in TABLE IV, incorporating $\mathcal{L}_{1}$ enhances emotion recognition accuracy by an average of 0.95% on downstream datasets compared to its absence.
III-C Discriminative Emotional Representation Learning
@@
According to psychological prior theories, the cognitive process of emotion is closely tied to the interaction between an individual and their environment. However, a natural semantic gap exists between these theories and the design of deep learning models. A significant challenge lies in designing effective fusion modules that integrate the feature representations of the full-scene image $I$ and the human position map $B$ to derive more nuanced and discriminative emotional representations. To address this challenge, we develop four methods, each targeting different degrees of this semantic gap.
@@
The full-scene image $I$ is processed by a YOLOv3 detector [39] and undergoes an image scaling transformation to obtain the human position map $B\in\mathbb{R}^{N\times H\times W\times C_{1}}$. The human position map $B$ is subjected to similar operations as $I$, passing through the ViT encoder and a linear layer to extract the CLS-token, denoted as $\beta\in\mathbb{R}^{N\times 1\times C_{3}}$. The feature representations of the body position map $\beta$ and the full-scene image $\alpha$ are reduced in dimensionality to $\mathbb{R}^{N\times C_{3}}$ and $\mathbb{R}^{N\times C_{3}}$, respectively.
To effectively fuse global scene features $\alpha\in\mathbb{R}^{N\times C}$ and human pose features $\beta\in\mathbb{R}^{N\times C}$, we propose a hierarchical fusion framework with four strategies ($\Gamma_{1}$$\Gamma_{4}$) tailored to varying degrees of semantic gaps in affective computing.
III-C1 Static Channel-weighted Fusion
To quantify the necessity of complex fusion strategies, we propose a parameter-free baseline fusion strategy, Static Channel-weighted Fusion, $\Gamma_{4}$, is defined as follows:
$\Upsilon=\lambda\odot\alpha+\beta,$ (2)
where $\odot$ represents the element-wise (Hadamard) product, and $\lambda\in\mathbb{R}^{C}$ is a predefined vector of weights. This method serves the following purposes:
  • Fixed Importance Distribution: Assume that the importance distribution between different semantic features ($\alpha$ and $\beta$) is fixed;
  • No Learnable Parameters: Provide a baseline fusion strategy with no learnable parameters, enabling direct evaluation of more complex fusion methods;
  • Semantic Gap Evaluation: Quantitatively assess the effectiveness of other strategies in reducing the semantic gap (see TABLE VI).
III-C2 Dynamic Weighted Fusion
To account for the dynamic variation in the importance of scene and human features (e.g., environment-dominant vs. human-dominant emotions), we propose a dynamic weighted fusion method, $\Gamma_{1}$, defined as follows
@@ III-C2 Dynamic Weighted Fusion
$\Upsilon=\operatorname{LayerNorm}\left(\sum_{i=1}^{\kappa}\left(\boldsymbol{\alpha}_{i}\odot\boldsymbol{\alpha}+\boldsymbol{\beta}_{i}\odot\boldsymbol{\beta}\right)\right)+\mathbf{b}.$ (3)
This method serves the following purposes:
  • Simpler Semantic Structures: Designed for datasets where scene clues are sparse, with either scene or human semantics dominating, leading to simpler semantic structures;
  • Learnable Weights: Employs multiple sets of learnable weights, $\lambda_{i},\mu_{i}$, to implement $\kappa$ predefined fusion modes, capturing a variety of semantic patterns;
  • Adaptive Emotional Mapping: Facilitates the model’s learning of adaptive emotional mapping strategies, such as when to prioritize body position or scene features;
  • Training Stability: Uses Layer Normalization (LayerNorm) to suppress outliers and improve training stability, particularly when there are significant scale differences between $\alpha$ and $\beta$ via:
    $\operatorname{LayerNorm}(\mathbf{x})=\frac{\mathbf{x}-\mathbb{E}[\mathbf{x}]}{\sqrt{\operatorname{Var}[\mathbf{x}]+\epsilon}}\odot\boldsymbol{\gamma}+\boldsymbol{\beta}.$ (4)
III-C3 Attention-driven Fusion
To handle fine-grained semantic ambiguities, we propose an attention-driven fusion strategy, $\Gamma_{2}$, defined as follows:
$\Upsilon=\operatorname{Swish}\left(\sum_{i=1}^{\kappa}\left(\operatorname{softmax}\left(\alpha\eta_{i}\right)\odot\alpha+\operatorname{softmax}\left(\beta\xi_{i}\right)\odot\beta\right)\right)+\epsilon.$ (5)
This strategy is primarily introduced to:
  • Complex Semantic Structures: Address datasets with rich scene clues and complex semantic structures;
  • Attention Mechanisms: Utilize multiple attention mechanisms, represented by $\eta_{i}$ and $\xi_{i}$, to generate $\kappa$ distinct attention distributions, allowing the model to adaptively select emotional mapping strategies based on sample characteristics.
III-C4 Gated Attention Fusion
To capture additional patterns and handle more complex scenarios, we propose a gated attention fusion method, $\Gamma_{3}$, defined as follows:
$\Delta_{i}=\operatorname{softmax}(\alpha\phi_{i})\odot\alpha+\operatorname{softmax}(\beta\varphi_{i})\odot\beta,$ (6a)
$\Delta_{con}=\operatorname{concat}(\Delta_{1},\Delta_{2},\cdots,\Delta_{\kappa}),$ (6b)
$\Delta_{m}=\operatorname{ReLU}(\Delta_{con}W_{1}+b_{1})W_{2}+b_{2},$ (6c)
$\Delta_{g}=\Delta_{m}\odot\sigma(\Delta_{m}W_{3}+b_{3}),$ (6d)
$\Upsilon=\operatorname{ReLU}(\Delta_{m}\odot\sigma(\Delta_{g}))+\epsilon.$ (6e)
The introduction of this method serves the following purposes:
  • High Noise and Occlusion Handling: Handle complex datasets with high noise and significant occlusion;
  • Complementary Alignment: Capture complementary alignment patterns between background and human features through multiple independent parameters, enhancing the model’s expressive power;
  • Noise Filtering: Employ a gating mechanism ($\sigma$) to dynamically filter noise and retain key cross-modal features.
III-D Distilling Rich Semantic Knowledge from CLIP
CLIP has demonstrated remarkable potential in cross-modal learning by bridging linguistic and visual modalities. Traditionally, language and vision processing have been treated as separate tasks; however, CLIP simultaneously models these modalities through contrastive learning, significantly improving multimodal alignment. Building on this success, we propose an emotion-aware pretraining framework that distills knowledge efficiently through two key strategies: Integration of vision-language contrastive learning paradigms during distillation, implemented via the proposed Contrastive Feature Alignment (CFA) Loss. Effective transfer of CLIP-acquired visual semantic knowledge to emotion-specific models, achieved through the Semantic-aware Visual Alignment (SVA) Loss.
We construct descriptive texts by naturally linking label attribute from the large-scale emotional dataset EmoSet, thereby creating numerous emotion image-text pairs. we input descriptive text $T$ and full-scene image $I$ into CLIP’s text encoder and image encoder, respectively, obtaining text features $\Lambda\in\mathbb{R}^{N\times C_{3}}$ and visual features $\Omega\in\mathbb{R}^{N\times C_{3}}$.
Given text features $\Lambda\in\mathbb{R}^{N\times C_{3}}$, CLIP visual features $\Omega\in\mathbb{R}^{N\times C_{3}}$, and visual emotional representations $\Upsilon\in\mathbb{R}^{N\times C_{3}}$, we first perform $\ell_{2}$-normalization:
$\Lambda^{\prime}_{i}=\frac{\Lambda_{i}}{\|\Lambda_{i}\|_{2}},\quad\Omega^{\prime}_{i}=\frac{\Omega_{i}}{\|\Omega_{i}\|_{2}},\quad\Upsilon^{\prime}_{i}=\frac{\Upsilon_{i}}{\|\Upsilon_{i}\|_{2}},$ (7)
where $\quad\forall i\in\{1,\ldots,N\}.$ This projects features onto a unit hypersphere, eliminating magnitude variations while preserving directional semantics.
Two cross-modal similarity matrices are computed as follows:
$A=\tau\cdot\Upsilon^{\prime}\Lambda^{\prime\top},\\ C=\tau\cdot\Omega^{\prime}\Lambda^{\prime\top}.$ (8)
where $\tau$ is a learnable temperature parameter.
The following content will present the core contributions of this subsection.
III-D1 Contrastive Feature Alignment Loss
To achieve efficient knowledge transfer through cross-modal contrastive alignment, we propose a Contrastive Feature Alignment (CFA) Loss $\mathcal{L}_{2}$, is defined as follows:
$\mathcal{L}_{2}=1-\frac{1}{N}\sum_{i=1}^{N}\left\langle\frac{A_{i}}{\|A_{i}\|_{2}},\frac{C_{i}}{\|C_{i}\|_{2}}\right\rangle,$ (9)